# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

### v4

@@ Abstract
Vision-language models (VLMs) have achieved impressive progress in natural image reasoning, yet their potential in medical imaging remains underexplored. Medical vision-language tasks demand precise understanding and clinically coherent answers, which are difficult to achieve due to the complexity of medical data and the scarcity of high-quality expert annotations. These challenges limit the effectiveness of conventional supervised fine-tuning (SFT) and Chain-of-Thought (CoT) strategies that work well in general domains. To address these challenges, we propose Med-R1, a reinforcement learning (RL)-enhanced vision-language model designed to improve generalization and reliability in medical reasoning. Built on the DeepSeek strategy, Med-R1 adopts Group Relative Policy Optimization (GRPO) to encourage reward-guided learning beyond static annotations. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities. Med-R1 achieves a 29.94% improvement in average accuracy over its base model Qwen2-VL-2B, and even outperforms Qwen2-VL-72B—a model with 36$\times$ more parameters. To assess cross-task generalization, we further evaluate Med-R1 on five question types. Med-R1 outperforms Qwen2-VL-2B by 32.06% in question-type generalization, also surpassing Qwen2-VL-72B. We further explore the thinking process in Med-R1, a crucial component for the success of Deepseek-R1. Our results show that omitting intermediate rationales (No-Thinking-Med-R1) not only improves in-domain and cross-domain generalization with less training, but also challenges the assumption that more reasoning always helps. These findings suggest that in medical VQA, it is not reasoning itself, but its quality and domain alignment, that determine effectiveness. Together, these results highlight that RL improves medical reasoning and generalization, enabling efficient and reliable VLMs for real-world deployment.
@@ 1 Introduction
To address these challenges, we propose Med-R1, a reinforcement learning (RL)-based framework for enhancing the generalizability and interpretability of medical VLMs. Unlike SFT, which aligns outputs to fixed supervision and often leads to shortcut learning, RL encourages exploration of diverse reasoning strategies through reward signals—without requiring explicit CoT annotations [11]. We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO [13] that stabilizes training via rule-based rewards and group-relative comparisons. These mechanisms reduce computational overhead while encouraging clinically grounded reasoning, making GRPO well-suited for medical tasks where scalability and reliability are essential.
@@
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
As shown in Figure 1, we evaluate Med-R1 across eight diverse medical imaging modalities, including CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy, and X-ray. These modalities cover a wide range of clinical imaging—from macroscopic anatomy to cellular-level and functional assessments. Med-R1 (2B parameters) achieves 69.91% average accuracy, a 29.94% gain over its base model Qwen2-VL-2B, and even outperforms the 72B-parameter Qwen2-VL-72B (Table 3), highlighting the benefit of RL-driven adaptation. We further assess generalizability across five question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 improves question-type generalization accuracy by 32.06% over Qwen2-VL-2B and surpasses Qwen2-VL-72B in this setting as well (Table 6). These demonstrate that RL enhances both parameter efficiency and task generalizability in medical VLMs.
Finally, inspired by recent findings on reasoning behavior in rule-based RL for classification tasks [14], we further investigate the role of intermediate rationales in Med-R1. A widely held assumption in VLM development is that generating step-by-step reasoning (“Think”) enhances generalization and performance. However, our empirical analysis reveals that this may not hold in domain-specific settings like medical VQA. We compare two RL post-training strategies: one that guides the model to produce explicit rationales before answering, and another that directly outputs the final answer without any intermediate reasoning steps. Surprisingly, we observe that the latter (No-Thinking) strategy often yields better generalization across both imaging modalities and question types. We hypothesize that in the absence of high-quality CoT supervision, RL-trained models fall back on their pre-trained reasoning patterns, which are typically tuned for general-domain tasks. Under large domain shifts—such as differences in image characteristics, task goals, and clinical language—this can result in hallucinated or misaligned rationales that impair performance. These findings suggest that it is not the presence of reasoning per se, but rather its quality and domain alignment, that determine its utility in medical vision-language reasoning.
@@
Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
@@
    Multi-modality multi-task reasoning medical VLM. We propose Med-R1, the first reasoning medical VLM capable of supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.) across five unique tasks. We demonstrate that RL-based fine-tuning effectively promotes modality-specific as well as cross-modality reasoning in the medical domain without the need for token-level supervision as in SFT. The proposed Med-R1 is capable of generating step-by-step, accurate, and clinically plausible explanations.
@@
    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating explicit thinking processes improves generalization. In medical VQA, we find that RL without reasoning steps yields more accurate predictions, while free-form reasoning—learned from general-domain data—can lead to hallucinations under domain shift. This highlights that in specialized domains—where image types, task goals, and language style differ significantly from general VLM training—the quality of reasoning, rather than its quantity, is critical for generalization.
@@ 2 Related Works
Medical Reasoning and Interpretability Interpretable reasoning is critical in medical AI, with recent work exploring CoT prompting [19] and program-guided logic [4]. However, such methods rely on costly expert annotations [20], limiting scalability in clinical domains. Reinforcement learning offers an alternative by enabling emergent reasoning without explicit supervision. Concurrently, MedVLM-R1 [21] also applies GRPO-based RL to radiology VQA, which is conducted on a relatively small dataset without generalization analysis, making its reasoning robustness challenging to assess. Nonetheless, it highlights the growing interest in RL for medical VLMs.
@@ 3 Method
We adopt rule-based RL to encourage multimodal reasoning and generalization in medical domains. Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [22] and visual navigation [11]. Building on these insights, we extend its application to the medical domain and systematically assess its effectiveness in this context. Specifically, we leverage the popular RL-based post-training method GRPO [23] to train a large base MLLM [24] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs. We introduce the details of the algorithm, our reward design, and data structure below.
@@ 3.1 Group Relative Policy Optimization (GRPO)
Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2-VL-2B.
RL fine-tuned VLM
Train Test CT MRI X-Ray US Der FP OCT Micro Overall
CT 98.30 74.43 80.61 51.93 64.17 65.39 70.52 66.13 71.44
MRI 62.33 99.29 73.75 53.18 75.04 64.57 75.59 66.31 71.26
X-Ray 79.67 68.29 95.85 49.32 70.75 76.87 70.28 67.75 72.35
US 55.88 63.96 70.53 99.86 64.78 63.02 63.92 71.53 69.19
Der 53.81 67.36 73.31 49.80 95.10 70.22 64.03 67.93 67.71
FP 54.98 65.95 75.79 49.08 71.36 92.44 65.33 66.40 67.67
OCT 64.48 70.80 77.09 52.65 72.89 73.41 99.17 66.85 72.17
Micro 60.97 63.25 74.92 52.56 64.55 63.66 66.86 93.51 67.54
Overall 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
@@
Table 2: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2-VL-2B.
No-Think RL fine-tuned VLM
Train Test CT MRI X-Ray US Der FP OCT Micro Overall
CT 99.54 82.18 80.00 46.24 72.13 72.50 76.06 70.36 74.88
MRI 79.27 99.72 74.74 42.86 79.94 73.50 84.67 69.10 75.48
X-Ray 78.37 69.94 97.09 46.43 70.75 80.97 69.81 70.63 73.00
US 59.82 66.53 73.37 100.00 70.60 68.03 68.63 79.28 73.28
Der 56.34 67.94 73.87 46.14 98.00 69.67 67.45 67.57 68.37
FP 59.55 68.62 73.93 51.59 72.43 93.99 72.64 70.36 70.39
OCT 80.47 73.50 77.15 49.08 65.08 70.95 100.00 67.21 72.93
Micro 64.15 66.56 76.84 53.13 68.91 70.95 73.47 98.56 71.50
Overall 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.48
@@ 3.2 No-Thinking-Med-R1
Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
@@
[…]
@@
Qwen2-VL-2B (GRPO + Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
Qwen2-VL-2B (GRPO + Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
Qwen2.5-VL-3B (GRPO + Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57

### v5

@@ Abstract
[…]
@@ 1 Introduction
To address these challenges, we propose Med-R1, a reinforcement learning (RL)-based framework for enhancing the generalizability and interpretability of medical VLMs. Unlike SFT, which aligns outputs to fixed supervision and often leads to shortcut learning, RL encourages exploration of diverse reasoning strategies through reward signals—without requiring explicit CoT annotations [11]. We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO (Proximal Policy Optimization) [13] that stabilizes training via rule-based rewards and group-relative comparisons. These mechanisms reduce computational overhead while encouraging clinically grounded reasoning, making GRPO well-suited for medical tasks where scalability and reliability are essential.
@@
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, OCT (Optical coherence tomography ), Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
As shown in Figure 1, we evaluate Med-R1 across eight diverse medical imaging modalities, including CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy, and X-ray. These modalities cover a wide range of clinical imaging—from macroscopic anatomy to cellular-level and functional assessments. Med-R1 (2B parameters) achieves 69.91% average accuracy, a 29.94% gain over its base model Qwen2-VL-2B, and even outperforms the 72B-parameter Qwen2-VL-72B (Table 4), highlighting the benefit of RL-driven adaptation. We further assess generalizability across five question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 improves question-type generalization accuracy by 32.06% over Qwen2-VL-2B and surpasses Qwen2-VL-72B in this setting as well (Table 8). These demonstrate that RL enhances both parameter efficiency and generalizability in medical VLMs.
Finally, we investigate the impact of intermediate rationales in Med-R1. Conventional wisdom suggests that explicit step-by-step reasoning (“Think”) enhances generalization, yet our results show this is not always true in medical VQA. We compare three RL post-training strategies: (1) Think, which generates rationales before answering; (2) No-Think, which directly predicts the answer; and (3) our proposed Think After, which performs concise reasoning after prediction. We find that No-Think improves generalization across modalities, while Think often causes hallucinated rationales due to domain shift. In contrast, Think After preserves interpretability without sacrificing accuracy, striking the best balance between reliability and explainability. These findings reveal that in specialized domains, the effectiveness of reasoning depends more on its quality, timing, and domain alignment than its length, challenging the notion that “more thinking is better.”
@@
Med-R1 a comprehensive and systematic study of rule-based RL for medical reasoning across eight imaging modalities (Figure 2). We summarized our contribution as follows:
@@
    Multi-modality and multi-task medical reasoning VLM. We propose Med-R1, a comprehensive and systematic study of rule-based reinforcement learning for medical reasoning, supporting eight imaging modalities (CT, MRI, Ultrasound, etc.) across five distinct clinical tasks. We demonstrate that RL-based fine-tuning effectively promotes modality-specific as well as cross-modality reasoning in the medical domain without the need for token-level supervision as in SFT. The proposed Med-R1 is capable of generating step-by-step, accurate, and plausible explanations.
@@
    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating longer or more explicit reasoning chains necessarily improves generalization. We find that reinforcement learning without explicit reasoning often yields higher accuracy, as free-form reasoning learned from general-domain data can induce hallucinations under domain shift. However, this may reduce reliability and interpretability in medical applications. To address this, we introduce Think-After—a reasoning scheme where the model provides rationalization for its chosen answer after the prediction. This design preserves interpretability while mitigating the instability introduced by lengthy reasoning chains, offering a balance between accuracy and explainability. Our findings suggest that the quality and timing of reasoning are critical for robust generalization.
@@ 2 Related Works
Medical Reasoning and Interpretability Interpretable reasoning is critical in medical AI, with recent work exploring CoT prompting [18] and program-guided logic [4]. However, such methods rely on costly expert annotations [19], limiting scalability in clinical domains. Reinforcement learning offers an alternative by enabling emergent reasoning without explicit supervision. Concurrently, MedVLM-R1 [20] also applies GRPO-based RL to radiology VQA, which is focuses on a single radiology-specific setting (training on MRI and testing on CT and X-ray) and reports generalization within radiology, using roughly 600 training samples. Nonetheless, it highlights the growing interest in RL for medical VLMs.
@@ 3 Method
We adopt rule-based RL to encourage multimodal reasoning and generalization in medical domains. Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [21] and visual navigation [11]. Building on these insights, we extend its application to the medical domain and systematically assess its effectiveness in this context. Specifically, we leverage the popular RL-based post-training method GRPO [22] to train a large base MLLM [23] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs. We introduce the details of the SFT and RL algorithm, our reward design, and data structure below.
Supervised Fine-tuning: We performed supervised fine-tuning (SFT) on the Qwen2-VL-2B and Qwen2.5-VL-3 models. Specifically, we fine-tuned the model using mixed-precision training (bfloat16) with gradient accumulation and checkpointing enabled for memory efficiency. The maximum sequence length was set to 4096 tokens, allowing the model to process full visual-text reasoning sequences. Training was conducted for one epoch with a learning rate of $2\times 10^{-5}$, cosine learning-rate scheduling, and a warm-up ratio of 0.1. The per-device batch size was 4, and the gradient accumulation step was 16, resulting in an effective batch size of 64.
@@ 3.1 Group Relative Policy Optimization (GRPO)
Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2.5-VL-3B.
RL fine-tuned Qwen2.5-VL-3B
Train Test CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
CT $94.35\pm 0.79$ $70.85\pm 1.10$ $86.50\pm 1.70$ $35.44\pm 2.05$ $66.77\pm 2.53$ $76.41\pm 2.51$ $84.08\pm 2.48$ $66.58\pm 2.79$ $72.62\pm 0.65$
MRI $72.79\pm 1.53$ $98.57\pm 0.29$ $72\pm 1.80 38.38$$38\pm 2.07$ $68.30\pm 2.53$ $79.78\pm 2.32$ $85.97\pm 2.36$ $67.12\pm 2.79$ $74.33\pm 0.64$
X-Ray $78.90\pm 1.39$ $59.73\pm 1.20$ $93.50\pm 1.21$ $38.28\pm 2.10$ $65.85\pm 2.57$ $76.05\pm 2.55$ $78.89\pm 2.77$ $65.14\pm 2.79$ $69.54\pm 0.68$
Ultrasound $62.20\pm 1.67$ $63.00\pm 1.19$ $79.57\pm 1.98$ $98.84\pm 0.46$ $64.09\pm 2.57$ $72.22\pm 2.60$ $78.54\pm 2.77$ $69.82\pm 2.75$ $73.53\pm 0.65$
Dermoscopy $64.30\pm 1.65$ $60.39\pm 1.18$ $80.56\pm 1.92$ $40.21\pm 2.12$ $84.99\pm 1.91$ $71.86\pm 2.69$ $74.76\pm 2.89$ $65.23\pm 2.75$ $67.79\pm 0.70$
Fundus $68.59\pm 1.59$ $61.32\pm 1.19$ $81.98\pm 1.86$ $37.42\pm 2.10$ $67.84\pm 2.49$ $90.62\pm 1.68$ $79.13\pm 2.71$ $66.76\pm 2.79$ $69.21\pm 0.67$
OCT $81.06\pm 1.34$ $73.92\pm 1.06$ $83.65\pm 1.80$ $35.78\pm 2.03$ $68.99\pm 2.53$ $80.87\pm 2.32$ $98.70\pm 0.77$ $67.12\pm 2.79$ $73.76\pm 0.65$
Microscopy $65.50\pm 1.62$ $62.34\pm 1.21$ $80.31\pm 1.95$ $37.37\pm 2.10$ $64.24\pm 2.60$ $1.40\pm 2.64 77$$7.36\pm 2.83$ $88.02\pm 1.94$ $68.32\pm 0.69$
Overall $73.46\pm 1.51$ $68.77\pm 1.12$ $83.72\pm 1.79$ $45.21\pm 2.14$ $68.88\pm 2.45$ $77.40\pm 2.50$ $82.18\pm 2.53$ $69.47\pm 2.74$ $71.14\pm 0.67$
@@
Table 2: Cross-Modality Generalization Performance of Med-R1 with No-think RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2.5-VL-3B
No-Think RL fine-tuned Qwen2.5-VL-3B
Train Test CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
CT $98.27\pm 0.45$ $74.77\pm 1.04$ $85.33\pm 1.70$ $37.61\pm 2.05$ $69.07\pm 2.53$ $78.87\pm 2.46$ $87.26\pm 2.30$ $67.75\pm 2.75$ $74.87\pm 0.64$
MRI $69.70\pm 1.56$ $99.61\pm 0.15$ $81.73\pm 1.89$ $34.91\pm 2.03$ $75.19\pm 2.34$ $80.51\pm 2.32$ $89.98\pm 2.00$ $67.48\pm 2.75$ $74.89\pm 0.65$
X-Ray $80.10\pm 1.36$ $69.78\pm 1.14$ $95.79\pm 0.99$ $34.33\pm 2.05$ $71.29\pm 2.41$ $81.79\pm 2.28$ $83.14\pm 2.54$ $68.65\pm 2.79$ $73.11\pm 0.63$
Ultrasound $56.40\pm 1.70$ $64.93\pm 1.18$ $80.25\pm 1.92$ $100.0\pm 0.00$ $68.38\pm 2.53$ $75.32\pm 2.60$ $76.30\pm 2.89$ $71.71\pm 2.66$ $74.16\pm 0.66$
Dermoscopy $55.45\pm 1.70$ $66.48\pm 1.15$ $80.37\pm 1.95$ $37.42\pm 2.12$ $92.88\pm 1.38$ $71.86\pm 2.69$ $76.53\pm 2.83$ $67.12\pm 2.70$ $68.51\pm 0.69$
Fundus $58.65\pm 1.68$ $66.08\pm 1.16$ $82.79\pm 1.83$ $34.33\pm 2.10$ $68.68\pm 2.49$ $93.81\pm 1.41$ $79.72\pm 2.71$ $65.86\pm 2.79$ $68.74\pm 0.69$
OCT $80.38\pm 1.36$ $75.02\pm 1.05$ $86.81\pm 1.67$ $36.11\pm 2.05$ $72.21\pm 2.41$ $81.60\pm 2.28$ $99.88\pm 0.18$ $65.32\pm 2.79$ $74.67\pm 0.64$
Microscopy $0.17\pm 1.73 66$$6.59\pm 1.16$ $83.10\pm 1.86$ $36.60\pm 2.10$ $1.29\pm 2.45 77$$7.05\pm 2.46$ $80.66\pm 2.71$ $97.66\pm 0.86$ $71.64\pm 0.67$
Overall $69.89\pm 1.59$ $72.91\pm 1.08$ $84.52\pm 1.73$ $43.91\pm 2.14$ $73.62\pm 2.37$ $80.10\pm 2.41$ $84.18\pm 2.47$ $71.44\pm 2.65$ $72.57\pm 0.66$
Table 3: Cross-Modality Generalization Performance of Med-R1 with Think-after RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2.5-VL-3B.
Think-after RL fine-tuned Qwen2.5-VL-3B
Train Test CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
CT $96.64\pm 0.62$ $72.53\pm 1.08$ $85.45\pm 1.67$ $36.31\pm 2.03$ $66.16\pm 2.53$ $79.78\pm 2.37$ $86.20\pm 2.30$ $68.20\pm 2.75$ $73.91\pm 0.65$
MRI $63.65\pm 1.67$ $98.87\pm 0.26$ $81.11\pm 1.92$ $33.70\pm 2.03$ $71.82\pm 2.45$ $81.51\pm 2.32$ $86.91\pm 2.30$ $67.57\pm 2.70$ $73.14\pm 0.67$
X-Ray $79.20\pm 1.42$ $72.70\pm 1.10$ $94.61\pm 1.08$ $31.58\pm 2.00$ $70.21\pm 2.45$ $80.60\pm 2.32$ $85.97\pm 2.30$ $68.65\pm 2.70$ $72.94\pm 0.66$
Ultrasound $59.06\pm 1.70$ $66.23\pm 1.15$ $80.25\pm 1.95$ $99.37\pm 0.34$ $67.38\pm 2.53$ $74.95\pm 2.55$ $78.77\pm 2.77$ $67.93\pm 2.75$ $74.24\pm 0.64$
Dermoscopy $56.77\pm 1.71$ $66.86\pm 1.15$ $80.37\pm 1.95$ $34.72\pm 2.07$ $89.51\pm 1.68$ $75.05\pm 2.55$ $78.66\pm 2.77$ $67.30\pm 2.75$ $68.65\pm 0.67$
Fundus $58.01\pm 1.68$ $65.89\pm 1.16$ $81.30\pm 1.86$ $32.84\pm 2.03$ $68.68\pm 2.53$ $92.53\pm 1.55$ $79.36\pm 2.77$ $66.22\pm 2.79$ $68.10\pm 0.69$
OCT $79.98\pm 1.37$ $75.45\pm 1.05$ $85.51\pm 1.73$ $34.14\pm 2.03$ $71.21\pm 2.45$ $81.69\pm 2.32$ $99.29\pm 0.53$ $66.94\pm 2.75$ $74.28\pm 0.65$
Microscopy $59.27\pm 1.68$ $64.22\pm 1.19$ $81.11\pm 1.89$ $34.47\pm 2.05$ $63.86\pm 2.60$ $73.68\pm 2.64$ $78.42\pm 2.77$ $86.67\pm 2.03$ $67.71\pm 0.69$
Overall $69.07\pm 1.62$ $72.84\pm 1.10$ $83.72\pm 1.76$ $42.14\pm 2.12$ $71.10\pm 2.53$ $79.97\pm 2.37$ $84.20\pm 2.47$ $69.93\pm 2.65$ $71.62\pm 0.66$
@@ 3.2 No-Thinking Med-R1
Methods Modality CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
@@
$\text{LLaMA Adapter v2}^{\dagger}$ [29] 21.41 26.63 46.44 34.05 51.76 50.74 33.00 38.66 37.83
$\text{MiniGPT-4}^{\dagger}$ [30] 22.81 27.48 38.30 25.50 40.25 38.33 31.40 36.23 32.54
InternVL2 [31] 40.20 58.10 57.90 49.10 51.90 53.20 59.10 64.00 54.19
@@
Qwen2-VL-2B (Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
Qwen2-VL-2B (Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
Qwen2.5-VL-3B (Think) 73.46 68.77 83.72 45.21 68.88 77.40 82.18 69.47 71.14
Qwen2.5-VL-3B (Think-after) 69.07 72.84 83.72 42.14 71.10 79.97 84.20 69.93 71.62
Qwen2.5-VL-3B (Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57
3.3 Think-after Med-R1
As said in the previous section, removing both reasoning and format supervision during RL could sometimes improve performance. However, as medical AI systems must not only achieve high accuracy but also provide reasoning that physicians can review and validate. To address this, we introduce a new reasoning protocol termed Think-After, in which the model first predicts the answer and then generates a post-hoc rationale explaining that decision. This design preserves interpretability while minimizing the instability introduced by lengthy reasoning chains. Concretely, the instruction prompt is revised as: {Question}. Output the single-letter choice (A, B, C, D,...) in <answer>...</answer> tags. Then provide the reasoning step by step to explain why you chose this answer. where {Question} is replaced by each specific VQA item. By separating answer prediction and rationale generation, we reduce interference between reasoning and decision processes, achieving balanced trade-off between accuracy and interpretability.